An SR 71 flew with a top speed of over 400km/h. How long did it take to fly from Washington to Moscow, a distance of 8300 km.


1
Expert's answer
2021-02-01T12:41:12-0500

We can find the time that an SR 71 takes to fly from Washington to Moscow from the formula:


"t=\dfrac{s}{v}=\dfrac{8300\ km}{400\ \dfrac{km}{h}}=20.75\ h."